OpenStudy (anonymous):

The area of a parallelogram is A = bh. Solve for h (height) in this formula.

OpenStudy (anonymous):

You have h defined in the equation, now you must isolate it.

HanAkoSolo (jamierox4ev3r):

well if a=bh, then you divide both sides by b, which gives you \(\large\frac{a}{b}=h\), or \(h=\large\frac{a}{b}\). Does that make sense? @mathematicshead ? Next problem I want you to try on your own. :)

OpenStudy (anonymous):

& hows that @Anickyan

HanAkoSolo (jamierox4ev3r):

well @Anickyan has stated, h is right there in the equation. it is not 2h, h^2, just h. all you had to do was get it by itself

OpenStudy (anonymous):

What Jamierox4ev3r said. Try to do it with this equation, just to practice a bit: \[2a + 3 = \frac{ h }{ b } - 5\] Solve for h

HanAkoSolo (jamierox4ev3r):

hello @mathematicshead ? what should you do to get the fraction \(\Large\frac{h}{b}\) by itself? Think now, and whatever you do must be done to both sides.

OpenStudy (anonymous):

i dont understand any of this at all, thats why im on here

OpenStudy (anonymous):

You should have listened in class. And if your teacher is assigning you this work, and haven't taught you it yet, you'd probably better talk to him/her.

HanAkoSolo (jamierox4ev3r):

well think about how you would solve a regular equation. When faced with an equation, just see if you can add or subtract the sides :) for example : x+2=8 -2 -2 ------- x=6

HanAkoSolo (jamierox4ev3r):

@mathematicshead watch my process carefully and see if you understand: \(2a+3=\large\frac{h}{b}-5\) +5 +5 ----------------- 2a+8=\(\frac{h}{b}\) multiply \(\large\frac{h}{b}\) by b and everything by b and you get : 2ab+8b=h

HanAkoSolo (jamierox4ev3r):

hey you guys there? O_O @mathematicshead and @Anickyan ?

OpenStudy (anonymous):

.................

OpenStudy (anonymous):

Yes. mathematicshead, this is really basic algebra. How old are you, and what grade are you in?

OpenStudy (anonymous):

@mathematicshead ^